OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (−1, −2); perpendicular to the line 2x + 7y + 2 = 0

OpenStudy (anonymous):

remember, we talked about "slope form" of the equation...? we need that one now. re-write your equation in slope form: 7y=-2x-2 y=-2/7 x -2/7 the coefficient in a front of x is a slope: m=-2/7

OpenStudy (anonymous):

now you need a line perpendicular to that one, which means the slope of YOUR line will be: m1= 7/2 (m=-1/m) now you know the slope of your desired line and your know the point that it has to go through, then: y-(-2)=7/2 (x-(-1)) y+2=7/2 (x+1) y=7/2 x +7/2 -2 y=7/2 x +3/2 this is the equation of your line